To make concrete mix Monica pours 34 pounds of cement 68 pounds of sand 14 pounds of small pebbles and 19 pounds of large pebbles into a large wheel barrel if she pours the mixture into 9 equal size bags how much will each bag weigh
1
Expert's answer
2019-08-09T05:19:57-0400

1) Calculate the mass of the mixture M, adding masses of all of the ingredients (cement, sand, small pebbles, large pebbles):

M = 34 + 68 + 14 + 19 = 135 pounds - total mass of the mixture.

2) We pour the same amount of mixture into each bag, so to find out the mass m of one bag we just divide the total mass M of the mixture by the number of bags:

m = M / 9 = 135/9 = 15 - mass of each bag.

Answer: each bag will weigh 15 pounds.